# Counting rationals and diophantine approximation in missing-digit Cantor sets

### v1

@@ 1. Introduction
In this paper, we study the distribution of rational numbers around missing-digit sets. Let $b\geqslant 2$ be an integer, and let $D$ be a non-empty subset of $\{0,1,\dots,b-1\}$. Let $K_{b,D}$ be the set of real numbers in $[0,1]$ having at least one base $b$ expansion with digits only from the set $D$. We call such a set $K_{b,D}$ a missing-digit set with base $b$ and digit set $D.$ For example, the set $K_{3,\{0,2\}}$ is the middle-third Cantor set.
@@
When $\mathrm{supp}(p):=\{j\in\{0,\dots,b-1\}:p_{j}\neq 0\}=D$, the measure $\nu_{p}$ is supported on $K_{b,D}$. The case when $p$ is the uniform probability vector on $D$, that is, it gives weight #$\#D^{-1}$ to each element of $D$, is of special importance to us. In this case, we call the associated measure the natural measure on $K_{b,D}$ and denote it by $\nu_{b,D}$.
@@
We use the Vinogradov notations ,$\ll$,$\gg$,$\asymp$ , ≫ , ≍, as well as the Bachmann–Landau notation $O(\cdot)$. Let $f,g$ be complex-valued functions. We write $f\ll g$ or $f=O(g)$ if $|f|\leqslant C|g|$ pointwise, for some constant $C>0$. We write $f\asymp g$ if $f\ll g$ and $g\ll f$.
@@ 1.1. Counting rational numbers in missing-digit sets
$T^{\kappa}\log T\ll\mathcal{N}_{K}(T)\ll T^{2\kappa}.$ (1.1)
@@
$\mathcal{N}_{K}(T)\ll T^{2{\kappa}-\varrho}.$ (1.2)
@@
$\nu(W_{K}({\alpha}))=0\qquad({\alpha}>2).$ (1.3)
@@
Let $K=K_{b,D}$ be such that $b\geqslant 5$ and #$\#D=b-1$ or $b=4$ and $D=\{1,2,3\}$ or $\{0,1,2\}$. Let $\nu$ be the natural measure on $K.$ Then there exists an effectively-computable $\varrho>0$ such that
@@
#$\#\{p/q\in K:q_{\mathrm{int}}\leqslant T\}\ll T^{{\kappa}+\varepsilon}.$
@@ 1.2. Extrinsic approximation
$\mathrm{dim}_{\mathrm{H}}(W_{\mathbb{R}}({\alpha})\cap K)=\max\left\{\frac{2}{{\alpha}}+{\kappa}-1,\frac{{\kappa}}{{\alpha}}\right\}.$ (1.4)
@@
$\mathrm{dim}_{\mathrm{H}}(\mathbf{VWA}\cap K)=\mathrm{dim}_{\mathrm{H}}(K).$ (1.5)
The last named author made progress towards these conjectures in [14, Theorem A and Remark A], showing that Conjecture 1.6 holds for missing-digit sets with base exceeding $10^{7}$, few missing digits and ${\alpha}$ being sufficiently close to $2$ We are able to extend this result to missing-digit sets that may not have large bases.
@@
Let $K=K_{b,D}$ with $b\geqslant 7$ and #$\#D=b-1$. Let $\kappa=\mathrm{dim}_{\mathrm{H}}(K)$. Then there is an effectively-computable $\varrho>0$ such that
@@ Acknowledgements
We thank Dmitry Kleinbock for a helpful suggestion. We thank Fredrik Johansson for helpful advice on numerical computations involving interval arithmetic.
@@ 2. Fourier $\ell^{1}$ dimension and applications
Let $b\geqslant 5$ and $D\subset\{0,1,\ldots,b-1\}$ with #$\#D=b-1$, or let $b=4$ and $D=\{0,1,2\}$ or $\{1,2,3\}$. Then $\hat{\kappa}_{1}(\nu_{b,D})>1/2$.
@@
Let $b\geqslant 7$ and $D\subset\{0,1,\ldots,b-1\}$ with #$\#D=b-1$. Then
@@
$D=\{kd+a:k=0,\ldots,l-1\}$
for some $a\in\mathbb{Z}_{\geqslant 0}$ and some $d,l\in\mathbb{Z}_{\geqslant 1}$. Assume $D$ is a proper subset of $\{0,\ldots,b-1\}$. Let $\nu$ be the natural missing-digit measure on $K_{b,D}$. Then
@@ 3.2. From Fourier $\ell^{1}$ dimension to counting rationals
[…]
@@
$\sum_{Q\leqslant q\leqslant 2Q}\nu(A(q,{\delta}))\ll{\delta}Q,$ (3.1)
@@
We may assume that $Q$ is large. Put $\varepsilon=\hat{\kappa}_{1}(\nu)-v>0$. Applying Parseval’s formula for a suitable test function approximating the characteristic function of $A(q,\delta)$, see [14, Theorem 4.1] for details, we have
@@ 4.1. Lower bound
For $N=0$, the claim follows from the fact that $|\hat{\nu}(y)|\leqslant 1$ for all $y$. We assume that $N>0$, and the claim holds for $N-1$. By (4.2), we have
@@ 4.2. Comparing two Riemann sums
    there is a constant $c_{0}\in(0,1/2]$ such that for all $x\in\mathbb{R}/\mathbb{Z}$, there are at most $b-2$ values of $a\in\{0,\ldots,b-1\}$ such that $\operatorname{dist}(x+a/b,Z)\leqslant c_{0}/b$,
@@
Item (1) holds because $p$ is a probability vector. Item (2) holds because, $g(x)=|Q(e(x))|$ for a polynomial $Q$ of degree at most $b-1$. For Item (3), since $g(0)=1$, we may by symmetry assume that $0<x\leqslant 1/(2b)$. Observe that $\operatorname{Im}(e(jx))>0$ for all $j\in\{1,\ldots,b-1\}$, hence
@@
unless $p_{0}=1$. In either case, we have $|g(x)|>0$ for $|x|\leqslant 1/(2b)$. The existence of $C_{0}$ follows from the smoothness of $g$ at 00, continuity, and compactness. We turn to Item (4). We first show that $g$ cannot vanish along an arithmetic progression of step $1/b$ and length $b-1$ in $\mathbb{R}/\mathbb{Z}$. Suppose for a contradiction that $g$ vanishes at $x+a/b$ for some $x\in\mathbb{R}/\mathbb{Z}$ and all $a\in\{1,\ldots,b-1\}$. Then
@@
is a polynomial of degree at most $b-1$ that vanishes at $e(x+a/b)$ for $a=1,\ldots,b-1$. It follows that
@@
for some $\widetilde{c}\in\mathbb{C}$. Since $p$ is a probability vector, this forces $e(-x)=1$ and $\widetilde{c}=1/b$. However, we explicitly excluded $p=(1/b,\ldots,1/b)$, so this is not possible. This contradiction proves our claim. The existence of $c_{0}$ follows again by continuity and compactness. Items (5) and (6) are immediate from the definitions.
@@
since $b^{-(L-1-j_{0})}$ is the length of the interval $J$ and $b^{-n}$ is the spacing between the points in question. We sum this up for all $z\in Z$ and $n=0,\ldots,L-1$ taking into account that $|Z|\leqslant b-1$ and that, for $n=L-1$, the number of $z\in Z$ such that there is $x\in\mathcal{X}$ with $\operatorname{dist}(b^{L-1}x,z)\leqslant c_{0}b^{-1}$ is at most $b-2$. (This is by Item (4) in Lemma 4.4.) We also use that $J\in\mathcal{J}^{(1)}$, hence the number of $n\in\{0,\ldots,L-1\}$ for which $\{x\in\mathcal{X}:\text{ $\operatorname{dist}(b^{n}x,z)<c_{0}b^{n-L}$}\}$ is non-empty is at most $L^{3/4}/(2b)\leqslant b^{j_{0}-1}/2$. We then get
[…]
@@ 4.3. Completing the proofs
for some constant $C$. This proves the claim. We observe that
@@ 5. Analytic bounds
Let $b\in\mathbb{Z}_{\geqslant 3}$, and suppose $D\subset\{0,1,\ldots,b-1\}$ with #$\#D=b-1$. Then
@@
$\hat{\kappa}_{1}(\nu)\geqslant\frac{-\log\max_{x}b^{-1}\sum_{i=0}^{b-1}g(x+i/b)}{\log b}.$ (5.1)
@@ 6. Numerical estimates
${\mathrm{Lip}}(F_{L})\leqslant b^{L}{\mathrm{Lip}}(S_{L})\leqslant b^{L}\sum_{j=0}^{L-1}2\pi(b-1)\cdot b^{j}=2\pi b^{L}(b^{L}-1).$ (6.2)
@@
Using the software SageMath [15], we verified (6.3) for the missing-digit measures $\nu=\nu_{b,\{0,\ldots,b-1\}\backslash\{a\}}$ for $b=4$ and $a=0$, and for
@@
with $\tau=1/2$, $L=2$ and $\delta=10^{-5}$, except that for $b=5$, $a=1$, we chose $L=4$ and $\delta=5\cdot 10^{-7}$. We note that the missing-digit measure $\nu_{b,\{0,\ldots,b-1\}\backslash\{b-1-a\}}$ is the image of $\nu_{b,\{0,\ldots,b-1\}\backslash\{a\}}$ under the map $x\mapsto 1-x$, which does not change Fourier $\ell^{1}$ dimension. For this reason, it is enough to do the calculations for $a\in\{0,\ldots,\lfloor(b-1)/2\rfloor\}$. This proves Proposition 2.4 for $b\leqslant 6$.
@@
$\min_{x\in A_{\delta,L}}F_{L}(x)>b^{(1-\tau)L}+\delta\cdot 2b^{L}(b^{L}-1)\pi,$ (6.5)

### v2

@@ 1. Introduction
In this paper, we study the distribution of rational numbers around missing-digit sets. Let $b\geqslant 2$ be an integer, and let $D$ be a subset of $\{0,1,\dots,b-1\}$ with $|D|\geqslant 2$. Let $K_{b,D}$ be the set of real numbers in $[0,1]$ having at least one base $b$ expansion with digits only from the set $D$. We call such a set $K_{b,D}$ a missing-digit set with base $b$ and digit set $D.$ For example, the set $K_{3,\{0,2\}}$ is the middle-third Cantor set.
@@
When $\mathrm{supp}(p):=\{j\in\{0,\dots,b-1\}:p_{j}\neq 0\}=D$, the measure $\nu_{p}$ is supported on $K_{b,D}$. The case when $p$ is the uniform probability vector on $D$, that is, it gives weight $D^{-1}$ to each element of $D$, is of special importance to us. In this case, we call the associated measure the natural measure on $K_{b,D}$ and denote it by $\nu_{b,D}$.
@@
We use the Vinogradov notations $\ll,\gg,\asymp$, as well as the Bachmann–Landau notation $O(\cdot)$. Let $f,g$ be complex-valued functions. We write $f\ll g$ or $f=O(g)$ if $|f|\leqslant C|g|$ pointwise, for some constant $C>0$. We write $f\asymp g$ if $f\ll g$ and $g\ll f$.
@@ 1.1. Counting rational numbers in missing-digit sets
(1.1) $T^{\kappa}\log T\ll\mathcal{N}_{K}(T)\ll T^{2\kappa}.$
@@
(1.2) $\mathcal{N}_{K}(T)\ll T^{2{\kappa}-\varrho}.$
@@
(1.3) $\nu(W_{K}({\alpha}))=0\qquad({\alpha}>2).$
@@
Let $K=K_{b,D}$ be such that $b\geqslant 5$ and #D=b1\#D=b-1 or $b=4$ and $D=\{1,2,3\}$ or $\{0,1,2\}$. Let $\nu$ be the natural measure on $K.$ Then there exists an effectively-computable $\varrho>0$ such that
@@
$\{p/q\in K:q_{\mathrm{int}}\leqslant T\}\ll T^{{\kappa}+\varepsilon}.$
@@ 1.2. Extrinsic approximation
(1.4) $\mathrm{dim}_{\mathrm{H}}(W_{\mathbb{R}}({\alpha})\cap K)=\max\left\{\frac{2}{{\alpha}}+{\kappa}-1,\frac{{\kappa}}{{\alpha}}\right\}.$
@@
(1.5) $\mathrm{dim}_{\mathrm{H}}(\mathbf{VWA}\cap K)=\mathrm{dim}_{\mathrm{H}}(K).$
The last named author made progress towards these conjectures in [14, Theorem A and Remark A], showing that Conjecture 1.6 holds for missing-digit sets with base exceeding $10^{7}$, few missing digits and $\alpha$} being sufficiently close to $2.$ We are able to extend this result to missing-digit sets that may not have large bases.
@@
Let $K=K_{b,D}$ with $b\geqslant 7$ and #D=b1\#D=b-1. Let $\kappa=\mathrm{dim}_{\mathrm{H}}(K)$. Then there is an effectively-computable $\varrho>0$ such that
@@ Acknowledgements
We thank Dmitry Kleinbock for a helpful suggestion. We thank Fredrik Johansson for helpful advice on numerical computations involving interval arithmetic. Finally, we thank an anonymous referee for a careful reading of the paper.
@@ 2. Fourier $\ell^{1}$ dimension and applications
Let $b\geqslant 5$ and $D\subset\{0,1,\ldots,b-1\}$ with #D=b1\#D=b-1, or let $b=4$ and $D=\{0,1,2\}$ or $\{1,2,3\}$. Then $\hat{\kappa}_{1}(\nu_{b,D})>1/2$.
@@
Let $b\geqslant 7$ and $D\subset\{0,1,\ldots,b-1\}$ with #D=b1\#D=b-1. Then
@@
$D=\{kd+a:k=0,\ldots,l-1\},$
for some integers $a\geqslant 0$, $d\geqslant 1$ and $l\geqslant 2$. Assume $D$ is a proper subset of $\{0,\ldots,b-1\}$. Let $\nu$ be the natural missing-digit measure on $K_{b,D}$. Then
@@ 3.2. From Fourier $\ell^{1}$ dimension to counting rationals
where $\|\cdot\|$ denotes distance to the nearest integer.
@@
(3.1) $\sum_{Q\leqslant q\leqslant 2Q}\nu(A(q,{\delta}))\ll{\delta}Q,$
@@
We may assume that $Q$ is large. Put $\varepsilon=\hat{\kappa}_{1}(\nu)-v>0$. For an upper bound, we may replace the indicator function of $A(q,\delta)$ by
$x\mapsto\sum_{a\in\mathbb{Z}}\phi\left(\frac{qx-a}{{\delta}}\right).$
Here $\phi:[0,\infty)$ is a Schwartz function such that $\phi\geqslant 1$ on [1,1][-1,1] and $\hat{\phi}$ is supported on [1,1][-1,1], a construction of which is given in [9, Example 3.2]. Applying Parseval’s formula — see [14, Theorem 4.1] for details — gives
@@ 4.1. Lower bound
For $N=0$, the claim follows from the fact that $|\hat{\nu}(y)|\leqslant 1$ for all $y$. We assume that $N>0$, and the claim holds for N1N-1. By (4.2), we have
@@ 4.2. Comparing two Riemann sums
    there is a constant $c_{0}\in(0,1/2]$ such that for all $x\in\mathbb{R}/\mathbb{Z}$, there are at most b2b-2 values of $a\in\{0,\ldots,b-1\}$ such that $\operatorname{dist}(x+a/b,Z)\leqslant c_{0}/b$,
@@
Item (1) holds because $p$ is a probability vector. Item (2) holds because, $g(x)=|Q(e(x))|$ for a polynomial $Q$ of degree at most b1b-1.
For Item (3), since $g(0)=1$, we may by symmetry assume that $0<x\leqslant 1/(2b)$. Observe that $\operatorname{Im}(e(jx))>0$ for all $j\in\{1,\ldots,b-1\}$, hence
@@
unless $p_{0}=1$. In either case, we have $|g(x)|>0$ for $|x|\leqslant 1/(2b)$. The existence of $C_{0}$ follows from the smoothness of $g$ at 0, continuity, and compactness.
We turn to Item (4). We first show that $g$ cannot vanish along an arithmetic progression of step $1/b$ and length b1b-1 in $\mathbb{R}/\mathbb{Z}$. Suppose for a contradiction that $g$ vanishes at $x+a/b$ for some $x\in\mathbb{R}/\mathbb{Z}$ and all $a\in\{1,\ldots,b-1\}$. Then
@@
is a polynomial of degree at most b1b-1 that vanishes at $e(x+a/b)$ for $a=1,\ldots,b-1$. It follows that
@@
for some $\widetilde{c}\in\mathbb{C}$. Since $p$ is a probability vector, this forces e(x)=1e(-x)=1 and $\widetilde{c}=1/b$. However, we explicitly excluded $p=(1/b,\ldots,1/b)$, so this is not possible. This contradiction proves our claim. The existence of $c_{0}$ follows again by continuity and compactness.
Items (5) and (6) are immediate from the definitions.
@@
since $b^{-(L-1-j_{0})}$ is the length of the interval $J$ and $b^{-n}$ is the spacing between the points in question.
We sum this up for all $z\in Z$ and $n=0,\ldots,L-1$ taking into account that $|Z|\leqslant b-1$ and that, for n=L1n=L-1, the number of $z\in Z$ such that there is $x\in\mathcal{X}$ with $\operatorname{dist}(b^{L-1}x,z)\leqslant c_{0}b^{-1}$ is at most b2b-2. (This is by Item (4) in Lemma 4.4.) We also use that $J\in\mathcal{J}^{(1)}$, hence the number of $n\in\{0,\ldots,L-1\}$ for which $\{x\in\mathcal{X}:\text{ $\operatorname{dist}(b^{n}x,z)<c_{0}b^{n-L}$}\}$ is non-empty is at most $L^{3/4}/(2b)\leqslant b^{j_{0}-1}/2$. We then get
$|\{x\in\mathcal{X}:$ $\text{ $\operatorname{dist}(b^{n}x,Z)<c_{0}b^{n-L}$ for some $0\leqslant n\leqslant L-1$}\}|$
@@ 4.3. Completing the proofs
for some constant $C$. This proves the claim.
We observe that
@@ 5. Analytic bounds
Let $b\in\mathbb{Z}_{\geqslant 3}$, and suppose $D\subset\{0,1,\ldots,b-1\}$ with #D=b1\#D=b-1. Then
@@
(5.1) $\hat{\kappa}_{1}(\nu)\geqslant\frac{-\log\max_{x}b^{-1}\sum_{i=0}^{b-1}g(x+i/b)}{\log b}.$
@@ 6. Numerical estimates
(6.2) ${\mathrm{Lip}}(F_{L})\leqslant b^{L}{\mathrm{Lip}}(S_{L})\leqslant b^{L}\sum_{j=0}^{L-1}2\pi(b-1)\cdot b^{j}=2\pi b^{L}(b^{L}-1).$
@@
[…]
@@
[…]
@@
(6.5) $\min_{x\in A_{\delta,L}}F_{L}(x)>b^{(1-\tau)L}+\delta\cdot 2b^{L}(b^{L}-1)\pi,$